Grade 10Inorganic Chemistry

An electron has kinetic energy 2.14 ×10^-22 joule. The de broglie wavelength will be nearly

Sanskriti nayak
8 Years agoGrade 10

1 Answer

Subrata Ghosh
8 Years ago
We know de Broglie Wave length=h/√2mE. =6.626*10^(-34)/ √2*9.11*10^(-31)*2.14*10^(-27) =1.061*10^-5metre